1/1

John Ellerton,  John Ellerton,

Du lässt den Tag, o Gott, nun enden The day Thou gavest, Lord, is ended

tradukita de Raymund Weber

Argiope-informo: La aspekto de tiu ĉi versio estas barita. Supozeble la aŭtoro decidis tion.

1. The day Thou gavest, Lord, is ended, The darkness falls at Thy behest; To Thee our morning hymns ascended, Thy praise shall sanctify our rest.

2. We thank Thee that Thy church, unsleeping, While earth rolls onward into light, Through all the world her watch is keeping, And rests not now by day or night.

3. As o’er each continent and island The dawn leads on another day, The voice of prayer is never silent, Nor dies the strain of praise away.

4. The sun that bids us rest is waking Our brethren ’neath the western sky, And hour by hour fresh lips are making Thy wondrous doings heard on high.

5. So be it, Lord; Thy throne shall never, Like earth’s proud empires, pass away: Thy kingdom stands, and grows forever, Till all Thy creatures own Thy sway.

Traduko de la Angla poemo “The day Thou gavest, Lord, is ended” de John Ellerton (∗1826-12-16 – †1893-06-15) en la Germanan de Raymund Weber.

Arg-130-256 (2008-11-08 15:30:25)

Diese Übertragung ins Deutsche ist abgedruckt in dem Gesangbuch der Katholischen Kirche in Deutschland “Gotteslob”, in der neueren Ausgabe. Der Text mit den Muziknoten findet sich auch auf der Heimseite http: // www. documentacatholicaomnia. u/ 01p/ 2006-09-10,_SS_Benedictus_ XVI,_Ves er_in_der_Kathedrale_in_ Muenchen,_GE. pdf .

Verkinto de tiu ĉi Angla poemo estas John Ellerton (∗1826-12-16 – †1893-06-15), publikigita 1870.

Arg-130-255 (2008-11-08 14:47:39)

Information about the author can be found on the website http: // www. cyberhymnal. org/ bio/ e/ l/ l/ ellerton_j. htm .

http://www.poezio.net/version?poem-id=130&version-id=256,255